OpenStudy (anonymous):

which ordered pair is the solution to the system of equations below? x+3y=7 x+2y=10

OpenStudy (shaik0124):

(16,-3) is the ordered pair of these equations

OpenStudy (anonymous):

can you explain how you got that please?

OpenStudy (anonymous):

I would use substitution.

OpenStudy (anonymous):

7-3y=x so 7-3y+2y=10

OpenStudy (anonymous):

7-y=10

OpenStudy (anonymous):

-3=y

OpenStudy (anonymous):

x+3(-3)=7

OpenStudy (anonymous):

x=-2

OpenStudy (shaik0124):

i solved by multiplying both equations by 3and first i got y value and substitute y vale in any equaton u will get x value

OpenStudy (anonymous):

You would multiply them both by 1/3

OpenStudy (anonymous):

... not 3.

OpenStudy (shaik0124):

the eqtns r in the form when u multi[ly by 3 on both sides 3x+9y=21 3x+6y=30 now subtract both equations then u will get 3y=-9 then y=-3 substitute y=-3 in any equation u get 16

OpenStudy (shaik0124):

multiplying 3 is correct@lord

OpenStudy (shaik0124):

hope the answer is done

OpenStudy (anonymous):

.so by multiplying by three the goal is to get rid of one of the variables right?

OpenStudy (shaik0124):

yes

OpenStudy (anonymous):

and u always subtract the equations?

OpenStudy (shaik0124):

when u get 1 value it will b easy to find other

OpenStudy (anonymous):

okay thanks alot.. i understand now :)

OpenStudy (shaik0124):

its way of saying thnks :)

OpenStudy (amistre64):

you can save a step by multiplying one of them by -3 instead, that preps it for subracting

OpenStudy (shaik0124):

u can do anyway yaar

OpenStudy (amistre64):

im not sure why the use of 3, did i read something in the post wrong?

OpenStudy (amistre64):

3 and -3 work fine, a little redundant, but fine ida gone with multiplying the second by -1

OpenStudy (amistre64):

x+3y=7 x+2y=10 ; *-1 x+3y = 7 -x -2y =-10 ; now combine ----------- y = -3

OpenStudy (shaik0124):

nyhow u can do

OpenStudy (anonymous):

soo if you multiply the bottom by a -1 then u dont have to multiply the top one by -1?

OpenStudy (shaik0124):

not required

OpenStudy (amistre64):

systems of equations have the property that scaling an equation doesnt change the variables outcomes

OpenStudy (amistre64):

the multiplying is just to change the form so that you can combine them to easily eliminate one of the variables from the process

OpenStudy (anonymous):

oh okay

OpenStudy (amistre64):

i simple example is of that property is: y = x has the same solutions as 2y=2x or y/3=x/3, or ny=nx

OpenStudy (anonymous):

okay so how do you know if the lines are parallel, or intersecting? does this apply to this type of problem or no?

OpenStudy (amistre64):

if one eqaution is a scalar of the other, they are the same line if one equation is almost a scalar of the other, and they differ only by a constant, they are parallel

OpenStudy (amistre64):

for example x+y = 3 , we can scale this by any factor and its still the same line 2x+2y = 6, or -5x-5y=-15 the line itself is not changing

OpenStudy (amistre64):

but, if they differ by only the constant term they are not the same line, but run in the same direction x+y = 3 2x+2y=3 would be parallel

OpenStudy (amistre64):

if they dont meet either of these criteria, the lines will intersect at some point

OpenStudy (anonymous):

so wouldnt this be identicle? or would it be parallel? y=-2+3 5y=-10x+15

OpenStudy (amistre64):

lets test it out, multiply the top by 5

OpenStudy (amistre64):

and your missing the "x" but i assume that a stypo :)

OpenStudy (anonymous):

oh yes its -2x sorry

OpenStudy (amistre64):

(y=-2x+3) *5 5y=-10x+15 5y= -10x+15 5y=-10x+15 same is same is same

OpenStudy (amistre64):

when lines are the same, they are called "coinciding" parallel refers to separate lines

OpenStudy (anonymous):

okay! sorry im asking soo many questions i just want to understand :) thanks alot!

OpenStudy (amistre64):

understanding is good ;)

OpenStudy (anonymous):

haha yeah